The method of standard additions was used to determine nitrite in a soil sample. A 1.00-ml portion of the sample was mixed wiBased on previous failed attempts, I think the answer is around 5.80x10^-4.
Answer #1

Let m moles of nitrite were present in the original sample.

Absorbance is given by beer lambert's law:

A=alc

0.3 = \varepsilon \: l x ( m / 23 )

\varepsilon \: l = 6.9 / m

Now,

Moles of nitrite from standard solution added = 10-3 x 10-3 = 10-6

Total moles of Nitrite after dilution = m + 10-6

A=alc

0.48 = ( 6.9 / m ) x ( m + 10-6 / 52)

3.6174 m =  m + 10-6

m = 3.82 x 10-7 moles

Concentration in original solution = 3.82 x 10-7 moles / 0.051 L

Concentration in original solution = 7.49 x 10-6 M